An 8.00-π‘˜π‘” point mass and a 12.0-π‘˜π‘” point mass are held in place 50.0 π‘π‘š apart. A particle of mass π‘š is released from a point between the two masses 20.0 π‘π‘š from the 8.00-π‘˜π‘” mass along the line connecting the two fixed masses. Find the magnitude and direction of the acceleration of the particle.Β 


1
Expert's answer
2021-12-20T10:26:39-0500
"a=(6.67\cdot10^{-11})\frac{8}{0.2^2}-(6.67\cdot10^{-11})\frac{12}{0.3^2}\\a=4.45\cdot10^{-9}\frac{m}{s^2}"

To the 8.00-π‘˜π‘” point mass.